# revtex4-2Repair the float

### v1

@@ IV.2 Multimodality
Next, we inspect the properties of the posterior distribution of coordinates in more detail. MCMC algorithms are computationally expensive, and one might wonder if we could do away with sampling altogether—be it HMC, BIGUE, or another algorithm—by quantifying uncertainty with Laplace’s approximation or a variational approach [75] instead. These approximations are attractive because they yield error bars for all quantities nearly for free, and they have strong theoretical backing, e.g., the Bernstein–von Mises theorem that guarantees convergence to the approximated distribution in the big data regime under certain conditions. Unfortunately, for small graphs and $\mathbb{S}^{1}$, at the very least, we find that they are not appropriate and that MCMC is truly necessary.
@@
Treating our sample as truly representative of the posterior distribution, we computed the p-values of the Henze-Zirkler normality test for all pairs of parameters and found the largest to be $4.82\times 10^{-14}$ for the example synthetic graph. This is not surprising when looking at the marginal distribution of the pair of coordinates shown in Fig. 5a: the bulk of the distribution is skewed, and there is a region of very low probability mass next to the bulk. All marginalizations of a multivariate normal yield another multivariate normal; thus, finding one non-normal marginal is enough to reject it as a model for the whole.
@@
It could be that only synthetic graphs show non-normal posterior distributions, but it turns out that the embedding of well-known graphs not explicitly generated by the model behaves the same way. For example, the posterior of Zachary’s Karate Club is also not normal, with a maximal p-value of $10^{-4}$ for the Henze-Zirkler normality test; Fig. 5b shows that joint marginals for this graph are not ellipsoids.
@@ IV.3 Observational study
To conclude our analysis, we embed a collection of empirical networks with BIGUE and contrast our results with Mercator’s; the experiments are summarized in Fig. 6 (see Appendix for a table giving numerical values in a table). Both algorithms reproduce the observed density and clustering (transitivity) of the graph. However, as we have argued above, Mercator yields a single embedding and thus ignores parameter uncertainty, leading to tighter estimated intervals.
@@
Properties that only depend on coordinates, like the global hierarchy level, have no associated ground truth in observational data and can only be compared across different embedding algorithms. Except for the macaque neural network, we find that Mercator systematically yields higher values of the hierarchy than BIGUE’s, usually by a large margin. This suggests that graphs are less hierarchical than they might seem if we only used Mercator embedding as the source of truth.
Finally, the link prediction results clearly show that finding a good representation for data is somewhat of an orthogonal goal to optimizing for prediction performance [76]. These results suggest that Mercator overfits to the observed graph to provide a better prediction of the links, while the posterior estimate selects the most plausible model while discounting statistical fluctuations.
V Conclusion
@@
Hyperbolic random graphs are unquestionably useful and can explain many properties observed in empirical networks, but existing estimation algorithms neglect the significant uncertainty and multimodality that can be present in these models, while off-the-shelf uncertainty quantification methods fall short. We provided evidence that the state-of-the-art dynamic HMC is unable to sample the posterior of a Bayesian hyperbolic embedding model and that supplementing random cluster transformations to a generic random walk is sufficient to sample embeddings of small graphs. Maximum likelihood estimators provided by Mercator overfit the coordinates do not provide error bars on embedding. This motivated us to introduce BIGUE, an MCMC algorithm with cluster transformations that provide an accurate posterior sample for the error bars on the graph and embedding properties.
As future work, we believe that generalizing cluster transformation on the $D+1$-sphere could be promising for developing effective sampling algorithms for the $\mathbb{H}^{D}$ hyperbolic model [49]. It’s very likely that HMC can be tweaked to perform better since there is a finite number of discontinuities, and this should also help a generalization in higher dimensions. Finally, the effectiveness of cluster transformations for sampling suggests that they could also be used to improve existing likelihood maximization algorithms.

### v2

@@ III.2 Multimodality
Next, we inspect the properties of the posterior distribution of coordinates in more detail. MCMC algorithms are computationally expensive, and one might wonder if we could do away with sampling altogether—be it HMC, BIGUE, or another algorithm—by quantifying uncertainty with Laplace’s approximation or a variational approach [87] instead. These approximations are attractive because they yield error bars for all quantities nearly for free, and they have strong theoretical backing, e.g., the Bernstein–von Mises theorem that guarantees convergence to the approximated distribution in the big data regime under certain conditions. Unfortunately, for small graphs and $\mathbb{S}^{1}$, at the very least, we find that the posterior distributions are not multivariate normal distributions and that MCMC is truly necessary.
@@
Treating our sample as truly representative of the posterior distribution, we computed the p-values of the Henze-Zirkler normality test [88] for all pairs of parameters. This test compares the empirical characteristic function to its pointwise limit under the null hypothesis that the data follows a multivariate normal distribution. We found the largest p-value for this test to be $1.11\times 10^{-25}$ for the example synthetic graph. This is not surprising when looking at the marginal distribution of the pair of coordinates shown in Fig. 6a: The bulk of the distribution is skewed, and there is a region of very low probability mass next to the bulk. All marginalizations of a multivariate normal yield another multivariate normal; thus, finding one non-normal marginal is enough to reject it as a model for the whole.
@@
It could be that only synthetic graphs show non-normal posterior distributions, but it turns out that the embedding of well-known graphs not explicitly generated by the model behaves the same way. For example, the posterior of Zachary’s Karate Club is also not normal: one pair of parameters has a p-value of $0.092$ for the Henze-Zirkler normality test, but the complete posterior has a p-value of 00 numerically; Fig. 6b shows that joint marginals for this graph are not ellipsoids.
@@ III.3 Observational study
To conclude our analysis, we embed a collection of empirical networks with BIGUE and contrast our results with Mercator’s; the experiments are summarized in Fig. 7 (see Appendix for a table giving numerical values). Both algorithms reproduce the observed density and clustering (transitivity) of the graph. The shortest paths average length is also reproduced by both algorithms except for the Zachary’s karate club and the Dutch literary critics networks, where both algorithms fail. In these cases, it is likely that the $\mathbb{S}^{1}$ model cannot offer a good representation altogether. And, as we have argued above, Mercator yields a single embedding and thus ignores parameter uncertainty, leading to tighter estimated intervals.
@@
Properties that only depend on coordinates, like the global hierarchy level, have no associated ground truth in observational data and can only be compared across different embedding algorithms. Except for the macaque neural network, we find that Mercator systematically yields higher values of the hierarchy than BIGUE’s, usually by a large margin. This suggests that graphs are less hierarchical than they might seem if we only had considered the Mercator embedding.
IV Conclusions
@@
Hyperbolic random graphs are unquestionably useful and can explain many properties observed in empirical networks, but existing estimation algorithms neglect the significant uncertainty and multimodality that can be present in these models, while off-the-shelf uncertainty quantification methods fall short. We provided evidence that the dynamic HMC is unable to sample the posterior of a Bayesian hyperbolic embedding model, and that supplementing random cluster transformations to a generic random walk is sufficient to sample embeddings of small graphs. While maximum likelihood estimators such as Mercator overfit the coordinates and do not provide error bars on embedding, our algorithm BIGUE provides an accurate posterior sample for the error bars on the graph and embedding properties.
As future work, we believe that generalizing cluster transformation on the $D+1$-sphere could be a promising avenue [52]. While each additional dimension would increase the computational cost by adding another parameter per vertex, this tradeoff may be worthwhile. The higher-dimensional space would allow vertices to navigate around each other more freely, potentially avoiding the barriers illustrated in Fig. 2.
Another challenge we leave for future work is to improve the algorithm’s overall computational efficiency. Currently, BIGUE’s main limitation is its poor scaling with vertex count—it struggles to properly sample graphs with more than 100 vertices in a reasonable timeframe. While basic optimizations like GPU-accelerated likelihood calculations and a more efficient programming language would help, the most promising speedups may lie in relying on HMC for the adjustment of positions. Naively using external implementations turned out to be too costly, but an integrated solution may work well. Other potential speed improvements include using negative sampling to approximate the likelihood [89] instead of computing the likelihood in $O(|V|^{2})$ operations; sampling vertices individually; incorporating graph symmetries during sampling—by swapping affected vertex positions—or even approximate symmetries [90].
Beyond sampling applications, the success of cluster transformations suggests they could enhance existing likelihood maximization algorithms like Mercator.
Data availability
The data were obtained from an online repository [91] or can be fetched from the original papers [79, 80, 81, 82, 83, 84, 85, 86]. All the data are also available with the code at https://doi.org/10.5281/zenodo.15272626.
Code availability
A Python implementation of BIGUE is available at https://doi.org/10.5281/zenodo.15189658. The Python code for the numerical analysis of this paper is available at https://doi.org/10.5281/zenodo.15272626.